OpenStudy (anonymous):

I have a fraction problem, I have the answer, but when I solve it, I get something totally different. 5/8x-1/3=2x-5/6

OpenStudy (anonymous):

is the x in 5/8x is in numerator or denominator

OpenStudy (anonymous):

I think it is just saying 5/8 of x. So I believe that it can be taken either or.

OpenStudy (anonymous):

multiply the whole equation with 24 15x-8=48x-20 33x=12 x=12/33=4/11